Kinkaid Co. is incorporated at the beginning of this year and engages in a number of transactions. The following journal entries impacted its stockholders’ equity during its first year of operations.

General Journal

Debit

Credit

a.

Cash

270,000

Common Stock, $25 Par Value

235,000

Paid-In Capital in Excess of Par Value, Common Stock

35,000

b.

Organization Expenses

200,000

Common Stock, $25 Par Value

127,000

Paid-In Capital in Excess of Par Value, Common Stock

73,000

c.

Cash

44,500

Accounts Receivable

15,500

Building

81,800

Notes Payable

59,800

Common Stock, $25 Par Value

52,000

Paid-In Capital in Excess of Par Value, Common Stock

30,000

d.

Cash

148,000

Common Stock, $25 Par Value

76,000

Paid-In Capital in Excess of Par Value, Common Stock

72,000


Required:
2. How many shares of common stock are outstanding at year-end?
3. What is the amount of minimum legal capital (based on par value) at year-end?
4. What is the total paid-in capital at year-end?
5. What is the book value per share of the common stock at year-end if total paid-in capital plus retained earnings equals $782,000?
Answer #1
2
Shares of common stock 19600 =(235000+127000+52000+76000)/25
3
Amount of minimum legal capital 490000 =19600*25
4
Total paid-in capital 700000 =490000+35000+73000+30000+72000
5
Book value per Common share
Choose Numerator / Choose Denominator = Book value per Common share
Stockholders' Equity applicable to Common shares / Number of common shares outstanding = Book value per Common share
782000 / 19600 = 39.90